# Leveraging LLMs, IDEs, and Semantic Embeddings for Automated Move Method Refactoring

### v1

@@ I Introduction
Second, MoveMethod refactoring requires global, project-level reasoning to determine the best target classes where to relocate a misplaced method. However, passing an entire project in the prompt is beyond the limited window size of current LLMs [27]. Even with larger window sizes, passing the whole project as context introduces noise and redundancy, as not all classes are relevant; instead this further distracts the LLM [27, 28]. We address the limited context size of LLMs by using retrieval augmented generation (RAG) to enhance the LLM’s input. Our two-step retrieval process combines mechanical feasibility (IDE-based static analysis) with semantic relevance (VoyageAI [29]), enabling our approach to make informed decisions and perform global project-level reasoning. We coin this approach refactoring-aware retrieval augmented generation, which addresses LLM hallucinations and context limitations while fulfilling the specific needs of MoveMethod refactoring (see Section III-C).
We designed, implemented, and evaluated these novel solutions as an IntelliJ IDEA plugin for Java, MM-assist (Move Method Professional). It synergistically combines the strengths of the LLM, IDE, static analysis, and semantic relevance. MM-assist generates candidates, filters LLM hallucinations, validates and ranks recommendations, and then finally executes the correct refactoring based on user approval using the IDE.
We designed a comprehensive, multi-methodology evaluation of MM-assist to corroborate, complement, and expand research findings: formative study, comparative study, replication of real-world refactorings, repository mining, user/case study, and questionnaire surveys. We compare MM-assist with the previous best in class approaches in their respective categories: JMove [10] – uses static analysis, feTruth [16] – uses Deep Learning, and HMove [30] – uses graph neural network to suggest moves and LLM to verify refactoring preconditions. These have been shown previously to outperform all previous MoveMethod recommendation tools. Using a synthetic corpus widely used by previous approaches, we found that MM-assist significantly outperforms them: for class instance methods, our Recall@1 and Recall@3 are 67% and 75%, respectively, which is an almost 1.75x improvement over previous state-of-the-art approaches (40% and 42%). Moreover, we extend the corpus used by previous researchers with 210 actual refactorings that we mined from OSS repositories in 2024 (thus avoiding LLM data contamination), containing both instance and static methods. We compared against JMove, HMove and feTruth on this real-world oracle, and found that MM-assist significantly outperforms them. Our Recall@3 is 80%, compared to 33% for the previous best tool, HMove– a 2.4x improvement. This shows that MM-assist’s recommendations better align with human developer best practices.
@@
  • Best Practices. We discovered a new set of best practices to overcome the LLM limitations when it comes to refactorings that require global reasoning. We automatically filter LLM hallucinations and conquer the LLM’s limited context size using refactoring-aware RAG.
@@ II Motivating Example
We illustrate the challenges of recommending MoveMethod using a real-world refactoring that occurred in the Elasticsearch project – a distributed open-source search and analytics engine. We illustrate the refactoring in Figure 1, and the full commit can be seen in  [31]. The resolvePolicy method (See 4 in Figure 1), originally part of the EsqlSession class, is misplaced. While EsqlSession handles parsing and executing queries, resolvePolicy is responsible for resolving and updating policies. Specifically, resolvePolicy accesses the field policyResolver (See 1) and parameters like groupedListener, policyNames, and resolution. Recognizing this misalignment, the developers refactored the code by moving resolvePolicy to the PolicyResolver class (not shown in the figure due to space constraints), updating the method body accordingly, and modifying the call sites (See 3). After the refactoring, both EsqlSession and PolicyResolver became more cohesive.
Automatically identifying such refactoring opportunities is essential for maintaining software quality, but it poses significant challenges for existing tools. We first applied HMove [30], the state-of-the-art MoveMethod technique. HMove, a classification tool, takes in $<$method, targetClass$>$ pairs, and gives a probability score indicating whether to move the method. After computing all 158 pairs of inputs, HMove executed for 1.5 hours and generated 36 MoveMethod recommendations. Its first and second highest recommendations to move are subfields and execute. Its third recommendation to move is resolvePolicy. However, HMove recommended moving resolvePolicy to the class FunctionRegistry, which is not an appropriate fit, as the method does not interact with it. This illustrates major shortcomings of classification-based tools like HMove: they need to be triggered on lots (avg. 145) of $<$method, targetClass$>$ pairs, which means long runtime. Furthermore, they overwhelm the users with so many recommendations to analyze, testing the developer’s patience and endurance. Moreover, they don’t provide actionable steps.
@@
After identifying that the method resolvePolicy is misplaced, a tool must find a suitable target class to move the method into. While the LLM was able to recommend the correct target class, it also responded with (i) two target classes (i.e., Resolution, ActionListener), which are plausible target classes, but are not the best-fit semantically for the method; (ii) two hallucinations, i.e., classes that do not exist (i.e., PolicyResolutionService, PolicyUtils as the LLM lacks project-wide context. A naive approach to address the LLM’s lack of project-wide understanding is to prompt it with the entire codebase. However, this is currently impractical due to the LLM’s context size limitations and inability to efficiently handle long contexts [33] (even though context limits continuously increase). Even state-of-the-art LLMs can’t process large projects like Elasticsearch in a single prompt without truncating crucial information. Moreover, the processing cost of large inputs with commercial LLM APIs is prohibitive.
These experiments reveal both the strengths and limitations of LLMs for MoveMethod refactoring. On the positive side, LLMs show proficiency in generating multiple suggestions and demonstrate an ability to identify methods that are semantically misplaced. However, they also exhibit significant limitations, including difficulty in suggesting appropriate target classes, and a high rate of irrelevant or infeasible suggestions. These limitations underscore the need for caution when relying on LLM-generated refactoring recommendations, and the need for a tedious manual analysis. Developers need to manually collect and re-analyze the suggestions, verify the suitability of each method for relocation, prompt the LLM again for suitable target classes, and meticulously identify and filter out hallucinations such as non-existent classes and methods that are impossible to move. In the example (Figure 1), a developer needs to sift through 5 candidate methods and, for each method, understand if any of the 5 or more proposed target classes are adequate. The developer analyzes 20+ $<$method, target class$>$ pairs before finding one they agree with.
@@ III Approach
Figure 2 shows the architecture and the steps performed by MM-assist. First, MM-assist applies a set of pre-conditions that filter out the methods that cannot be safely moved, such as constructors ( 1 in Figure 2). It then leverages vector embeddings from Language Models to identify methods that are the least cohesive with their host class ( 2 in Figure 2). In Figure 1, by comparing the embeddings of resolvePolicy and EsqlSession using cosine similarity, MM-assist detects that this method might be misplaced (§$\S$ III-B). Next, MM-assist passes the remaining candidates to the LLM (i.e., the method signature and the class body), which analyzes their relationships with the host class to prioritize the most promising MoveMethod recommendations ( 3 in Figure 2).
Once MM-assist identifies candidate methods, it systematically evaluates potential target classes from the project codebase. For the resolvePolicy method, which utilizes the enrichPolicyResolver field ( 1 in Figure 1), MM-assist initially identifies several candidate classes, including EnrichPolicyResolver and PolicyManager.
To narrow down the target classes, MM-assist calculates relevance scores between the candidate method and each potential target class – establishing a ranking ( 4 in Figure 2). We label this process as “refactoring-aware RAG”. Then, we feed the LLM with the narrowed-down list of target classes, and ask it to pick the best one ( 5 in Figure 2). In this case, it correctly selects EnrichPolicyResolver as the appropriate destination for resolvePolicy, aligning with the developers’ actual refactoring decision (see 3 in Figure 1). Finally, refactoring suggestions ( 6 in Figure 2) are presented to the user, and MM-assist leverages the IDE’s refactoring APIs to safely execute the chosen one automatically.
@@ III-A Important Concepts
[…]
@@
(Valid Refactoring Recommendations) These are recommendations that do not break the code. They are mechanically feasible: syntactically correct and successfully pass the preconditions as checked by the IDE. We differentiate between moving an instance and a static method:
@@ III-B Identifying Which Method To Move
Identify Least Compatible Methods via Embedding-Based Analysis. To further refine candidate methods, we use an embedding-based analysis. An embedding is a vector representation that captures the semantic features of an entity (methods and classes) based on their content and relationships. We leverage VoyageAI embeddings [29], specifically trained on code, as they more effectively capture the semantic relevance of programming constructs. We use VoyageAI [29] due to its state-of-the-art performance in code-related tasks. Vectors are generated for two inputs: one for the method body and another for the host class, excluding the method body. Excluding the method ensures that the class embedding remains unbiased by the method itself. We then calculate the cosine similarity between these vectors to assess how well each method semantically aligns with its host class.
Prioritize Methods with LLM Guidance. To make the analysis computationally tractable and fit the LLM context size, we further narrow candidate methods based on LLM’s understanding of class level design. We use the LLM to rank the existing methods in our suggestion pool. Using Chain-of-Thought (CoT) reasoning, we prompt the LLM to perform a structured analysis: evaluate each method’s purpose, cohesion and dependencies, summarize the host class’s responsibilities, and assess overall alignment of the method & class.
@@ III-C Recommending Suitable Target Classes
Target Class Retrieval - Instance Methods. As instance methods can be moved to a few suitable classes - we select the method’s parameter types and the host class’ field types as potential destinations. Then, we utilize the IDE’s preconditions to retain only valid MoveMethod suggestions.
Target Class Retrieval - Static Methods. We identify potential target classes within the project based on two key aspects: package proximity and utility class identification. Package proximity quantifies structural closeness in the package hierarchy by computing shared package segments (e.g., for org.example.core and org.example.utils, "org" and "example" are shared) normalized by the host package depth, providing an initial structural filter. Utility classes, identified through conventional naming patterns (containing "util" or "utility"), are prioritized as common targets for static methods. These heuristics are efficient filters to narrow down the search space of potential target classes. We rank potential target classes based on the above heuristics, with greater importance given to proximity. Our ranking function is:
@@
[…]
@@
Semantic Relevance-Based Target Class Ranking. While static analysis offers foundational understanding of valid refactoring opportunities, it often yields a broad set of potential target classes, as it lacks the ability to capture deeper semantic relationships. To augment the results of static analysis, we incorporate a semantic relevance analysis, which compares both the content and intent of the candidate method and target classes. To do this, we utilize VoyageAI’s vector embeddings [29] to compute the cosine similarity between the method body and potential target classes. We sort the target classes by their cosine similarity scores in descending order to select the most semantically relevant candidates for the LLM to analyze. To stay within the LLM’s context window, we limit the candidates to those fitting within a 7K token budget – typically accommodating 10-12 class summaries with their signatures.
Ranking Target Classes Using LLM. In the final phase, MM-assist asks the LLM for the best-suited target class, utilizing its vast training knowledge. To avoid context overflow, we create a concise representation of each target class, including its name, field declarations, DocString, and method signatures. The LLM then takes as input the method to be moved along with these summarized target class representations, returning a prioritized list of target classes.
@@ IV Empirical Evaluation
    How effective are LLMs at suggesting MoveMethod refactoring opportunities? This question assesses Vanilla LLMs’ ability to recommend MoveMethod through a formative study examining the quality of LLM suggestions.
@@
    How effective is MM-assist at suggesting MoveMethod refactoring opportunities? We evaluate the performance of MM-assist against the state-of-the-art tools, feTruth [16] and HMove [30] (representatives for DL approaches), and JMove [10] (representative for static analysis approaches). We used both a synthetic corpus used by other researchers and a new dataset of real refactorings from open-source developers.
@@ IV-A Subject Systems
[…]
@@
Real-world corpus. As refactorings in the real-world are often complex and messy [36], we decided to complement the synthetic dataset with a corpus of actual MoveMethod refactorings that open-source developers performed on their projects. This dataset allows us to determine whether various tools can match the rationale of expert developers in real-world situations. We construct this oracle using RefactoringMiner [37], the state-of-the-art for detecting refactorings [38].
@@
We took extra precautions to prevent LLM data contamination, ensuring that the LLMs used by MM-assist had no prior exposure to the data we tested and could not rely on previously memorized results. With GPT-4’s knowledge cutoff in October 2023, we focused our analysis on repository commits from January 2024 onward. We ran RefactoringMiner on the 25 most actively maintained Java repositories on GitHub, ranked by commit history, and with over 1000 stars.
@@
However, many of the MoveMethod reported by RefactoringMiner are false positives, often resulting from residual effects of other refactorings such as MoveClass (where an entire class is relocated to another package) or Extract Class (where a class is split into two, creating a new class along with the original). To filter out these false positives, we employed several techniques: first, we verified that both the source and target classes existed in both versions of the code (i.e., at the commit head and its previous head). We removed test methods, getters, setters, and overriden methods (they violate preconditions for a MoveMethod). For instance methods, we then checked if the method was moved to a field in the source class, to a parameter type, or if the signature of the moved method contained a reference to the source class. Starting from the instances detected by RefactoringMiner, we curated a dataset of 210 verified MoveMethod, with 102 static methods and 108 instance methods – on 12 popular open-source projects. On average, each project contains 8743 classes and 66306 methods spanning 1032344 LOCs. This oracle enables an evaluation of MM-assist on authentic refactorings made by experienced developers.

### v2

# Together We Are Better: LLM, IDE and Semantic Embedding to Assist Move Method Refactoring

@@ I Introduction
Second, MoveMethod refactoring requires global, project-level reasoning to determine the best target classes where to relocate a misplaced method. However, passing an entire project in the prompt is beyond the limited context window of current LLMs [27]. Even with larger window sizes, passing the whole project as context introduces noise and redundancy, as not all classes are relevant; instead this further distracts the LLM [27, 28]. We address the limited context window of LLMs by using retrieval augmented generation (RAG) to enhance the LLM’s input. Our two-step retrieval process combines mechanical feasibility (IDE-based static analysis) with semantic relevance (VoyageAI [29]), enabling our approach to make informed decisions and perform global project-level reasoning. We coin this approach refactoring-aware retrieval augmented generation, which addresses LLM hallucinations and context limitations while fulfilling the specific needs of MoveMethod refactoring (see Section III-C).
We designed, implemented, and evaluated these novel solutions as an IntelliJ IDEA plugin for Java, MM-assist (Move Method Assist). It synergistically combines the strengths of the LLM, IDE, static analysis, and semantic relevance. MM-assist generates candidates, filters LLM hallucinations, validates and ranks recommendations, and finally executes the correct refactoring based on user approval using the IDE.
We designed a comprehensive, multi-methodology evaluation of MM-assist to corroborate, complement, and expand research findings: formative study, comparative study, replication of real-world refactorings, repository mining, user/case study, and questionnaire surveys. We compare MM-assist with the previous best in class approaches in their respective categories: JMove [10] – uses static analysis, feTruth [16] – uses Deep Learning, and HMove [30] – uses graph neural network to suggest moves and LLM to verify refactoring preconditions. These have been shown to outperform all previous MoveMethod recommendation tools. Using a synthetic corpus widely used by previous approaches, we found that our tool significantly outperforms them: for class instance methods, our Recall@1 and Recall@3 are 67% and 75%, respectively, which is almost 1.75x improvement over previous state-of-the-art approaches (40% and 42%). Moreover, we extend the corpus used by previous researchers with 210 actual refactorings that we mined from OSS repositories in 2024 (thus avoiding LLM data contamination), containing both instance and static methods. We compared against JMove, HMove and feTruth on this real-world oracle, and found that MM-assist significantly outperforms them. Our Recall@3 is 80%, compared to 33% for the best baseline, HMove, a 2.4x improvement. This shows that MM-assist’s recommendations better align with developer best practices.
@@
  • Best Practices. We discovered a new set of best practices to overcome the LLM limitations when it comes to refactorings that require global reasoning. We automatically filter LLM hallucinations and conquer the LLM’s limited context window size using refactoring-aware RAG.
@@
  • Replication. We freely release MM-assist, the datasets we used in the experiments, LLM prompts, demo, etc, so others can build upon these [31].
@@ II Motivating Example
We illustrate the challenges of recommending MoveMethod using a real-world refactoring that occurred in the Elasticsearch project – a distributed open-source search and analytics engine. We illustrate the refactoring in Figure˜1, and the full commit can be seen in  [32]. The resolvePolicy method (See 4 in Figure˜1), originally part of the EsqlSession class, is misplaced. While EsqlSession handles parsing and executing queries, resolvePolicy is responsible for resolving and updating policies. Specifically, resolvePolicy accesses the field policyResolver (See 1) and parameters like groupedListener, policyNames, and resolution. Recognizing this misalignment, the developers refactored the code by moving resolvePolicy to the PolicyResolver class (not shown in the figure due to space constraints), updating the method body accordingly, and modifying the call sites (See 3). After the refactoring, both EsqlSession and PolicyResolver became more cohesive.
Automatically identifying such refactoring opportunities is essential for maintaining software quality, but it poses significant challenges for existing tools. We first applied HMove [30], the state-of-the-art MoveMethod technique. HMove, a classification tool, takes in <<method, targetClass>> pairs, and gives a probability score indicating whether to move the method. After computing all 158 pairs of inputs, HMove executed for 1.5 hours and generated 36 MoveMethod recommendations. Its first and second highest recommendations to move are subfields and execute. Its third recommendation to move is resolvePolicy. However, HMove recommended moving resolvePolicy to the class FunctionRegistry, which is not an appropriate fit, as the method does not interact with it. This illustrates major shortcomings of classification-based tools like HMove: they need to be triggered on lots (avg. 145) of <<method, targetClass>> pairs, which means long runtime. Furthermore, they overwhelm the users with so many recommendations to analyze, testing the developer’s patience and endurance. Moreover, they don’t provide actionable steps.
@@
After identifying that the method resolvePolicy is misplaced, a tool must find a suitable target class to move the method into. While the LLM was able to recommend the correct target class, it also responded with (i) two target classes (i.e., Resolution, ActionListener), which are plausible target classes, but are not the best-fit semantically for the method; (ii) two hallucinations, i.e., classes that do not exist (i.e., PolicyResolutionService, PolicyUtils) as the LLM lacks project-wide context. A naive approach to address the LLM’s lack of project-wide understanding is to prompt it with the entire codebase. However, this is currently impractical due to the LLM’s context size limitations and inability to efficiently handle long contexts [34] (even though context limits continuously increase). Even state-of-the-art LLMs can’t process large projects like Elasticsearch in a single prompt without truncating crucial information. Moreover, the processing cost of large inputs with commercial LLM APIs is prohibitive.
These experiments reveal both the strengths and limitations of LLMs for MoveMethod refactoring. On the positive side, LLMs show proficiency in generating multiple suggestions and demonstrate an ability to identify methods that are semantically misplaced. However, they also exhibit significant limitations, including difficulty in suggesting appropriate target classes, and a high rate of irrelevant or infeasible suggestions. These limitations underscore the need for caution when relying on LLM-generated refactoring recommendations, and the need for a tedious manual analysis. Developers need to manually collect and re-analyze the suggestions, verify the suitability of each method for relocation, prompt the LLM again for suitable target classes, and meticulously identify and filter out hallucinations such as non-existent classes and methods that are impossible to move. In the example (Figure˜1), a developer needs to sift through 5 candidate methods and, for each method, understand if any of the 5 or more proposed target classes are adequate. The developer analyzes 20+ <<method, target class>> pairs before finding one they agree with.
@@ III Approach
Figure˜2 shows the architecture and the steps performed by MM-assist. First, MM-assist applies a set of pre-conditions that filter out the methods that cannot be safely moved, such as constructors ( 1 in Figure˜2). It then leverages vector embeddings from Language Models to identify methods that are the least cohesive with their host class ( 2 in Figure˜2). In Figure˜1, by comparing the embeddings of resolvePolicy and EsqlSession using cosine similarity, MM-assist detects that this method might be misplaced ($\S$ III-B). Next, MM-assist passes the remaining candidates to the LLM (i.e., the method signature and the class body), which analyzes their relationships with the host class to prioritize the most promising MoveMethod recommendations ( 3 in Figure˜2).
Once MM-assist identifies candidate methods, it systematically evaluates potential target classes from the project codebase. For the resolvePolicy method, which utilizes the enrichPolicyResolver field ( 1 in Figure˜1), MM-assist initially identifies several candidate classes, including EnrichPolicyResolver and PolicyManager.
To narrow down the target classes, MM-assist calculates relevance scores between the candidate method and each potential target class – establishing a ranking ( 4 in Figure˜2). We label this process as “refactoring-aware RAG”. Then, we feed the LLM with the narrowed-down list of target classes, and ask it to pick the best one ( 5 in Figure˜2). In this case, it correctly selects EnrichPolicyResolver as the appropriate destination for resolvePolicy, aligning with the developers’ actual refactoring decision (see 3 in Figure˜1). Finally, refactoring suggestions ( 6 in Figure˜2) are presented to the user, and MM-assist leverages the IDE’s refactoring APIs to safely execute the chosen one automatically.
@@ III-A Important Concepts
(MoveMethod Refactoring) A MoveMethod refactoring moves a method m from a host class H (where it currently resides) to a target class T. We define a MoveMethod $\omega$" as a triplet $({\texttt{\footnotesize{m}}},{\texttt{\footnotesize{H}}},{\texttt{\footnotesize{T}}})$.
@@
(Valid Refactoring Recommendations) These are recommendations that do not break the code. They are mechanically feasible: they successfully pass the preconditions as checked by the IDE, thus resulting in syntactically and semantically equivalent code. We differentiate between moving an instance and a static method:
@@ III-B Identifying Which Method To Move
Identifying Semantically Mismatched Methods. For the remaining candidates, the next step is to find methods that are out of place in their current class. To automate this, we use code embeddings (VoyageAI [29]) to capture the semantic meaning of the code. An embedding is a vector representation that captures the semantic features of an entity (methods and classes) based on its content and relationships. We use VoyageAI due to its state-of-the-art performance in code-related tasks. We generate vectors for two inputs: one for the method body and another for the host class, excluding the method body. Excluding the method ensures that the class embedding remains unbiased by the method itself. We then calculate the cosine similarity between these vectors to assess how well each method semantically aligns with its host class.
Prioritizing Candidates with LLM-driven Reasoning. Finally, after identifying a pool of semantically misplaced methods, we use an LLM to rank them based on a broader perspective (i.e., class level design). We use the LLM to rank the existing methods in our suggestion pool. Using Chain-of-Thought (CoT) reasoning, we prompt the LLM to perform a structured analysis: evaluate each method’s purpose, cohesion and dependencies, summarize the host class’s responsibilities, and assess overall alignment of the method & class.
@@ III-C Recommending Suitable Target Classes
III-C1 Candidate Filtering by Method Type
The initial filtering process differs based on whether the method is an instance or static type, as they possess different relocation constraints.
Target Class Retrieval for Instance Methods
As instance methods can be moved to a few suitable classes, we select the method’s parameter types and the host class’ field types as potential destinations. Then, we utilize the IDE’s preconditions to retain only valid MoveMethod suggestions.
Target Class Retrieval for Static Methods
We identify potential target classes within the project based on two key aspects: package proximity and utility class identification. Package proximity quantifies structural closeness in the package hierarchy by computing shared package segments (e.g., for org.example.core and org.example.utils, "org" and "example" are shared) normalized by the host package depth, providing an initial structural filter. Utility classes, identified through conventional naming patterns (containing "util" or "utility"), are prioritized as common targets for static methods. These heuristics are efficient filters to narrow down the search space of potential target classes. We rank potential target classes based on the above heuristics, with greater importance given to proximity. Our ranking function is:
@@ III-C1 Candidate Filtering by Method Type
  • $\text{proximity}({\texttt{\footnotesize{T}}},{\texttt{\footnotesize{H}}})$ evaluates the package proximity between class T and the host class H
  • $\text{isUtility}({\texttt{\footnotesize{T}}})$ is a boolean function that returns 1 if T is a utility class, and 0 otherwise
@@
III-C2 Semantic Relevance-Based Target Class Ranking
While static analysis offers foundational understanding of valid refactoring opportunities, it often yields a broad set of potential target classes, as it lacks the ability to capture deeper semantic relationships. To augment the results of static analysis, we incorporate a semantic relevance analysis, which compares both the content and intent of the candidate method and target classes. To do this, we use VoyageAI’s vector embeddings [29] to compute the cosine similarity between the method body and potential target classes. We sort the target classes by their cosine similarity scores in descending order to select the most semantically relevant candidates for the LLM to analyze. To stay within the LLM’s context window, we limit the candidates to those fitting within a 7K token budget – typically accommodating 10-12 class summaries with their signatures.
III-C3 Ranking Target Classes Using LLM
In the final phase, MM-assist asks the LLM for the best-suited target class, utilizing its vast training knowledge. To avoid context overflow, we create a concise representation of each target class, including its name, field declarations, DocString, and method signatures. The LLM then takes as input the method to be moved along with these summarized target class representations, returning a prioritized list of target classes.
@@ IV Empirical Evaluation
    How effective are LLMs at suggesting opportunities for MoveMethod refactoring? This question assesses Vanilla LLMs’ ability to recommend MoveMethod in a formative study examining the quality of LLM suggestions.
@@
    How effective is MM-assist at suggesting opportunities for MoveMethod refactoring? We evaluate the performance of MM-assist against the state-of-the-art tools, feTruth [16] and HMove [30] (representatives for DL approaches), and JMove [10] (representative for static analysis approaches). We used both a synthetic corpus used by other researchers and a new dataset of real refactorings from open-source developers.
@@ IV-A Subject Systems
To evaluate LLMs’ capability when suggesting MoveMethod refactoring opportunities, we employed two distinct datasets: a synthetic corpus widely used by previous researchers [10, 36, 13, 30] and a new corpus that contains real-world refactorings that open-source developers performed. Each corpus comes along with a “gold set” $G$ of MoveMethod refactorings that a recommendation tool must attempt to match. We define $G$ as a set of MoveMethod refactorings (see  Definition˜III.1) - each containing a triplet of method-to-move, host class, and target class $({\texttt{\footnotesize{m}}},{\texttt{\footnotesize{H}}},{\texttt{\footnotesize{T}}})$.
IV-A1 Synthetic corpus
The synthetic corpus was created by Terra et al. [10] moving different methods m out of their original/host class H to a random destination class T. The researchers then created the gold set as tuples $({\texttt{\footnotesize{m}}},{\texttt{\footnotesize{H}}},{\texttt{\footnotesize{T}}})$, i.e., methods m that a tool should now move from H back to its original class T. This dataset moves only instance methods; it does not move static methods. This corpus consists of 10 open-source projects.
IV-A2 Real-world corpus
@@
As refactorings in the real-world are often complex and messy [37], we decided to complement the synthetic dataset with a corpus of actual MoveMethod refactorings that open-source developers performed on their projects. This dataset allows us to determine whether various tools can match the rationale of expert developers in real-world situations. We construct this oracle using RefactoringMiner [38], the state-of-the-art for detecting refactorings [39].
@@
We took extra precautions to prevent LLM data contamination, ensuring that the LLMs used by MM-assist had no prior exposure to the data we tested and could not rely on previously memorized results. With GPT-4’s knowledge cutoff in October 2023, we focused our analysis on repository commits from January 2024 onward.
@@
However, many of the MoveMethod reported by RefactoringMiner are not pure MoveMethod refactorings, often resulting from residual effects of other refactorings such as MoveClass (where an entire class is relocated to another package) or Extract Class (where a class is split into two, creating a new class along with the original)  [40]. These impure MoveMethod operations require additional changes to the code’s structure, to facilitate them (e.g. creating a new class, moving fields, etc.). To deal with the limited number of data points from MoveMethod refactorings detected by RefactoringMiner, we leveraged ‘Extract and Move Method’ refactoring detected by RefactoringMiner (a composite refactoring in which developers first extract a new method and then move it to a different class). Empirically, we found that detected ‘Extract and Move Method’s resulted in a greater number of pure refactorings than detected MoveMethod. This can be attributed to the reason that small sections of a methods are out of place, rather than the entire method, and performing the extract highlights the need for move. Moreover, the ‘Extract and Move Method’ refactoring is frequently performed in practice. We constructed our dataset by executing the extract method (within the IDE), and generating an intermediate version of the code for tools to analyze and generate recommendations. Then, we performed further analysis to validate that the target method can in-fact be moved to the target-class. To achieve this, we applied a process similar to prior work [16, 30]: first, we verified that both the source and target classes existed in both versions of the code (i.e., at the commit head and its previous head). We removed test methods, getters, setters, and overriden methods (they violate preconditions for a MoveMethod). For instance methods, we then checked if the method was moved to a field in the source class, to a parameter type, or if the signature of the moved method contained a reference to the source class. Starting from the instances detected by RefactoringMiner, we curated a dataset of 210 verified MoveMethod, with 102 static methods and 108 instance methods – on 12 popular open-source projects. On average, each project contains 8743 classes and 66306 methods spanning 1032344 LOCs. This oracle enables an evaluation of MM-assist on authentic refactorings made by experienced developers.